# Distribution estimation via Flow Matching with Lipschitz guarantees

### v1

@@ 7.2 Proofs of Section˜4
$\int_{t^{*}}^{1}|(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}})\frac{t^{\gamma}}{\sigma_{t}^{2}}\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{2}O\big{(}\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{\frac{1}{\kappa}}\big{)}|\;\mathrm{d}t$ $\leq\int_{t^{*}}^{1}|(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}})||\frac{1}{t^{\gamma}}O\big{(}\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{\frac{1}{\kappa}}\big{)}|\;\mathrm{d}t$
$\lesssim\int_{t^{*}}^{1}\gamma t^{-1-\frac{\kappa}{\gamma}}\sigma_{t}^{\frac{1}{\kappa}}\;\mathrm{d}t+\int_{t^{*}}^{1}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\frac{\sigma_{t}^{\frac{1}{\kappa}}}{t^{\frac{\gamma}{\kappa}}}\;\mathrm{d}t$
$\leq\gamma\int_{t^{*}}^{1}t^{-1-\frac{\kappa}{\gamma}}\;\mathrm{d}t+(t^{*})^{-\frac{\gamma}{\kappa}}\int_{t^{*}}^{1}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\sigma_{t}^{\frac{1}{\kappa}}\;\mathrm{d}t$
$=\frac{\gamma^{2}}{\kappa}((t^{*})^{-\frac{\kappa}{\gamma}}-1)+(t^{*})^{-\frac{\gamma}{\kappa}}\kappa(\sigma_{\min}^{\frac{1}{\kappa}}-\sigma_{t^{*}}^{\frac{1}{\kappa}}).$
Using that $\sigma_{\min}^{\frac{1}{\kappa}}-\sigma_{t^{*}}^{\frac{1}{\kappa}}\leq 1$ yields the following bound on the integral over the Lipschitz constant
$\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t\lesssim\vartheta^{2}(1+\log(\vartheta))+(t^{*})^{-2\gamma}(\gamma+1)+\frac{\gamma^{2}}{\kappa}((t^{*})^{-\frac{\kappa}{\gamma}}-1)+(t^{*})^{-\frac{\gamma}{\kappa}}\kappa.$
Proof of Theorem˜4.6.
For the proof we need the following additional definition:
@@
From (Boucheron et al., 2013, Theorem 3.20) we know that $\mathcal{N}(0,I_{d})$ satisfies the Poincaré inequality with Poincaré constant $1.$ Thus, via defining $g(x)\coloneqq f(\sigma x+\mu)$ and
$\operatorname{Var}(g(X))\leq\sigma^{2}\mathbb{E}[|\nabla g(X)|^{2}],$
we see that the Poincaré constant of $\mathcal{N}(\mu,\sigma^{2}I_{d})$ is $\sigma^{-2}$. Now we are ready to proof Theorem˜4.6.
Proof.
Property (III):
@@
$q\propto\exp\big{(}-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}\big{)}=\exp\Big{(}-\frac{1}{2}\big{(}1+\frac{t^{2}}{\sigma_{t}^{2}}\big{)}\big{(}|y|^{2}-\langle\frac{t}{\sigma_{t}^{2}}\frac{x}{1+\frac{t^{2}}{\sigma_{t}^{2}}},y\rangle\big{)}\Big{)}$ (20)
is a Gaussian distribution with variance $\big{(}1+\frac{t^{2}}{\sigma_{t}^{2}}\big{)}^{-1}I_{d}$. Hence the density defined above satisfies the Poincaré inequality with constant $\big{(}1+\frac{t^{2}}{\sigma_{t}^{2}}\big{)}$. Using the Holley-Strooke perturbation principle Holley and Stroock (1987), in the form of (Ledoux, 2001, Lemma 1.2), we can bound the Poincaré constant $\rho$ of the perturbed Gaussian via
$\rho\geq e^{-4L}\big{(}1+\frac{t^{2}}{\sigma_{t}^{2}}\big{)}.$
Thus
@@
we conclude that for all $t\in[0,1]$
$\operatorname{Cov}(Y^{x,t}_{i},(Y^{x,t}_{j})\leq e^{4L}.$
Hence we can set $C=e^{4L}$.
Property (II):
@@
$\mathbb{E}[\big{(}Y^{x,t}_{i}\big{)}^{2}]$ $=\frac{\int\big{(}\frac{x_{i}-z_{i}\sigma_{t}}{t}\big{)}^{2}\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}{\int\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}$
$=\big{(}\frac{x_{i}}{t}\big{)}^{2}-\frac{2x_{i}\sigma_{t}}{t^{2}}\frac{\int z_{i}\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}{\int\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}+\big{(}\frac{\sigma_{t}}{t}\big{)}^{2}\frac{\int z_{i}^{2}\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}{\int\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}.$
@@
$\mathbb{E}[(Y^{x,t}_{i})^{2}]-\mathbb{E}[Y^{x,t}_{i}]^{2}$ $=\big{(}\frac{x_{i}}{t}\big{)}^{2}-\frac{2x_{i}\sigma_{t}}{t^{2}}A(z_{i})+\big{(}\frac{\sigma_{t}}{t}\big{)}^{2}A(z_{i}^{2})-\big{(}\frac{x_{i}}{t}-\frac{\sigma_{t}}{t}A(z_{i})\big{)}^{2}$
$=\big{(}\frac{\sigma_{t}}{t}\big{)}^{2}\big{(}A(z_{i}^{2})-A(z_{i})^{2}\big{)}.$ (21)
@@
$p_{Z}(z)$ $=\frac{\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})}{\int\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}z}$
$=\frac{\exp\big{(}-\frac{|z|^{2}}{2}-\frac{|x-z\sigma_{t}|^{2}}{2t^{2}}-a(\frac{x-z\sigma_{t}}{t})\big{)}}{\int\exp\big{(}-\frac{|z|^{2}}{2}-\frac{|x-z\sigma_{t}|^{2}}{2t^{2}}-a(\frac{x-z\sigma_{t}}{t})\big{)}\;\mathrm{d}z}$
@@
$\mathbb{E}_{p_{Z}}[\partial_{z_{i}}\log(p(Z))]$ $=\int\partial_{z_{i}}p(z)\;\mathrm{d}z$
$=\int...\int_{z_{i}\in\mathbb{R}}\partial_{z_{i}}p(z)\;\mathrm{d}z_{i}\;\mathrm{d}(z_{1},...,z_{i-1},z_{z+1},...,z_{d})$
$=0.$
@@
$0=-\mathbb{E}\Big{[}\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\big{(}Z_{i}-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\big{)}-\frac{\sigma_{t}}{t}\partial_{z_{i}}a(\frac{x-Z\sigma_{t}}{t})\Big{]}$
$\Longleftrightarrow\quad$ $\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\big{(}\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\big{)}=\mathbb{E}\Big{[}\frac{\sigma_{t}}{t}\partial_{z_{i}}a(\frac{x-Z\sigma_{t}}{t})\Big{]}$
@@
where $L$ is from (14). From (22) we can see that this bounds the influence of the perturbation function $a$ on the expected value.
Now we bound the variance. To do so, we first note that
@@
$-\mathbb{E}_{p_{Z}}[\partial^{2}_{z_{i}}\log(p_{Z}(Z))]$ $=\big{(}1+(\frac{\sigma_{t}}{t})^{2}\big{)}+(\frac{\sigma_{t}}{t})^{2}\mathbb{E}_{p_{Z}}\big{[}\partial^{2}_{z_{i}}a(\frac{x-\sigma_{t}z}{t})\big{]}.$
@@
$\mathbb{E}_{p_{Z}}[(\partial_{z_{i}}\log(p(Z)))^{2}]=\mathbb{E}_{p_{Z}}\Big{[}\Big{(}\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}(Z_{i}-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})})+\frac{\sigma_{t}}{t}\partial_{z_{i}}a(\frac{x-\sigma_{t}Z}{t})\Big{)}^{2}\Big{]}$
$=\mathbb{E}_{p_{Z}}\Big{[}\Big{(}\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}(Z_{i}-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})})\Big{)}^{2}\Big{]}+2\mathbb{E}_{p_{Z}}\Big{[}\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}(Z_{i}-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})})\frac{\sigma_{t}}{t}\partial_{z_{i}}a(\frac{x-\sigma_{t}Z}{t})\Big{]}$
@@
$\mathbb{E}_{p_{Z}}$ $\Big{[}\Big{(}\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}(Z_{i}-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})})\Big{)}^{2}\Big{]}=\mathbb{E}_{p_{Z}}\Big{[}\Big{(}\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}(Z_{i}-\mathbb{E}[Z_{i}]+\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})})\Big{)}^{2}\Big{]}$
$=\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}^{2}\mathbb{E}_{p_{Z}}\Big{[}\Big{(}(Z_{i}-\mathbb{E}[Z_{i}])\Big{)}^{2}\Big{]}+2\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}^{2}(\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})})\mathbb{E}_{p_{Z}}\Big{[}(Z_{i}-\mathbb{E}[Z_{i}])\Big{]}$
$\quad+\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}^{2}\Big{(}\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big{)}^{2}$
$=\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}^{2}\Big{(}\mathbb{E}_{p_{Z}}\Big{[}\Big{(}(Z_{i}-\mathbb{E}[Z_{i}])\Big{)}^{2}\Big{]}+\Big{(}\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big{)}^{2}\Big{)}.$
@@
$\mathbb{E}_{p_{Z}}\Big{[}\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}(Z_{i}-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})})\frac{\sigma_{t}}{t}\partial_{z_{i}}a(\frac{x-\sigma_{t}Z}{t})\Big{]}$
$=\mathbb{E}_{p_{Z}}\Big{[}\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\big{(}Z_{i}-\mathbb{E}[Z_{i}]+\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\big{)}\frac{\sigma_{t}}{t}\partial_{z_{i}}a(\frac{x-\sigma_{t}Z}{t})\Big{]}$
$=\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\frac{\sigma_{t}}{t}\mathbb{E}_{p_{Z}}\Big{[}(Z_{i}-\mathbb{E}[Z_{i}])\partial_{z_{i}}a(\frac{x-\sigma_{t}Z}{t})\Big{]}$
$\quad+\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\frac{\sigma_{t}}{t}\big{(}\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\big{)}\mathbb{E}_{p_{Z}}\Big{[}\partial_{z_{i}}a(\frac{x-\sigma_{t}Z}{t})\Big{]}.$

### v2

@@ 6.1 Proofs of Section˜3
$\int_{t^{*}}^{1}\Big|\Big(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\Big)\frac{t^{\gamma}}{\sigma_{t}^{2}}\Big(\frac{\sigma_{t}}{t^{\gamma}}\Big)^{2}O\Big(\Big(\frac{\sigma_{t}}{t^{\gamma}}\Big)^{\frac{1}{\kappa}}\Big)\Big|\;\mathrm{d}t\leq\int_{t^{*}}^{1}\Big|\Big(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\Big)\Big|\Big|\frac{1}{t^{\gamma}}O\Big(\Big(\frac{\sigma_{t}}{t^{\gamma}}\Big)^{\frac{1}{\kappa}}\Big)\Big|\;\mathrm{d}t$
$\lesssim\int_{t^{*}}^{1}\gamma t^{-1-\frac{\gamma}{\kappa}}\sigma_{t}^{\frac{1}{\kappa}}\;\mathrm{d}t-\int_{t^{*}}^{1}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\frac{\sigma_{t}^{\frac{1}{\kappa}}}{t^{\frac{\gamma}{\kappa}}}\;\mathrm{d}t\leq\gamma\int_{t^{*}}^{1}t^{-1-\frac{\gamma}{\kappa}}\;\mathrm{d}t-(t^{*})^{-\frac{\gamma}{\kappa}}\int_{t^{*}}^{1}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\sigma_{t}^{\frac{1}{\kappa}}\;\mathrm{d}t.$
Bounding the last term with $\kappa((t^{*})^{-\frac{\gamma}{\kappa}}-1)+(t^{*})^{-\frac{\gamma}{\kappa}}\kappa\Big(\sigma_{1}^{\frac{1}{\kappa}}-\sigma_{t^{*}}^{\frac{1}{\kappa}}\Big)$ and using that $\sigma_{1}^{\frac{1}{\kappa}}-\sigma_{t^{*}}^{\frac{1}{\kappa}}\leq 1$ yields the following bound on the integral over the Lipschitz constant
$\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t\lesssim\vartheta^{2}(1+\log(\vartheta))+(t^{*})^{-2\gamma}(\gamma+1)+\kappa((t^{*})^{-\frac{\gamma}{\kappa}}-1)+(t^{*})^{-\frac{\gamma}{\kappa}}\kappa.\qed$
Preparations for the proof of Proposition˜3.6.
For the proof we need the following additional definitions:
@@
From (Boucheron et al., 2013, Theorem 3.20) we know that $\mathcal{N}(0,I_{d})$ satisfies the Poincaré inequality with Poincaré constant $1.$ Thus, via defining $g(x)\coloneqq f(\sigma x+\mu)$ and $\operatorname{Var}(g(X))\leq\sigma^{2}\mathbb{E}[|\nabla g(X)|^{2}],$ we see that the Poincaré constant of $\mathcal{N}(\mu,\sigma^{2}I_{d})$ is $\sigma^{-2}$.
Definition 6.4.
The standard Gaussian satisfies the $\log$-Sobolev inequality with $\log$-Sobolev constant $\lambda=1$ (Boucheron et al., 2013, Theorem 5.4). Analogously to Poincaré inequality, its easy to see that $\mathcal{N}(\mu,\sigma^{2}I_{d})$ satisfies the $\log$-Sobolev inequality with $\log$-Sobolev constant $\lambda=\sigma^{-2}$.
Now we are ready to prove Proposition˜3.6.
Proof of Proposition˜3.6.
Property (III), $\mathbb{P}^{*}$ of form (12):
We start by defining $\Phi(y)=\frac{|x-t^{\gamma}y|^{2}}{2\sigma_{t}^{2}}+V(y)$, such that $q(y)\propto\exp(-\Phi(y))$. For every $t$, we have that $\nabla^{2}\Phi(y)=\frac{t^{\gamma}}{\sigma_{t}^{2}}I+\nabla^{2}V(y)\gtrsim\nabla^{2}V(y)\gtrsim mI$. Thus $\mathbb{P}^{*}$ is strongly $\log$-concave. By Brascamp and Lieb (1976), we can bound
$\operatorname{Cov}(Y^{x,t})_{ij}\leq\sqrt{\operatorname{Var}(Y^{x,t}_{i})}\sqrt{\operatorname{Var}(Y^{x,t}_{j})}\leq\frac{1}{m}.$
Property (I) and (II), $\mathbb{P}^{*}$ of form (12):
We choose $t^{*}$ such that $\frac{\sigma_{t}^{2}}{t^{2\gamma}}L\leq\min(\frac{1}{2},\frac{\sigma_{t}}{t^{\gamma}})$. Our goal is to derive a representation of $\Sigma$, the covariance matrix of $q$. Using the divergence theorem, as common in multivariate Stein’s identity type results relating to Stein (1972), see e.g. Oates et al. (2017, Section 2.2.2), we obtain for the smooth function $f(y)=y-\mathbb{E}[Y]$ and $q\propto\exp(-\Psi)$
$\mathbb{E}[\nabla f(Y)]=\mathbb{E}[f(Y)(\nabla\Psi(Y))^{\top}]\quad\Longleftrightarrow\quad I=\mathbb{E}[(Y-\mathbb{E}[Y])(\nabla\Psi(Y))^{\top}].$
For $t>t^{*}>0$, we define $\mu\coloneqq\frac{x}{t^{\gamma}}$ and $s=\frac{\sigma_{t}}{t^{\gamma}}$ such that $\Psi(y)=\frac{|y-\mu|^{2}}{2s^{2}}+V(y)$. Then since $\mathbb{E}[Y-\mathbb{E}[Y]]=0$
$\mathbb{E}[(Y-\mathbb{E}[Y])(\nabla\Psi(Y))^{\top}]=\mathbb{E}\Big[(Y-\mathbb{E}[Y])\Big(\frac{Y-\mu}{s^{2}}+\nabla V(Y)\Big)^{\top}\Big]=\frac{1}{s^{2}}\Sigma+\mathbb{E}\big[(Y-\mathbb{E}[Y])(\nabla V(Y))^{\top}\big].$
Thus
$\Sigma=s^{2}I-s^{2}\mathbb{E}\big[(Y-\mathbb{E}[Y])(\nabla V(Y))^{\top}\big].$
A multivariate Taylor expansion yields
$\nabla V(Y)=\nabla V(\mathbb{E}[Y])+\nabla^{2}V(\mathbb{E}[Y])(Y-\mathbb{E}[Y])+R(Y),$
where $R$ depends on the bound of the third partial derivatives and $|R(Y)|\lesssim|Y-\mathbb{E}{Y}|^{2}$. Again, since $\mathbb{E}[Y-\mathbb{E}[Y]]=0$, we obtain
$\Sigma=s^{2}I-s^{2}\Sigma\nabla^{2}(V(\mathbb{E}[Y]))^{\top}-R^{\prime},$
where $|R^{\prime}|\lesssim\mathbb{E}[|Y-\mathbb{E}[Y]|^{3}]$. By Villani (2008, Theorem 21.2, Remark 12.4), who refers to Bakry and Émery (1985), which is in French, a strongly $\log$-concave function satisfies the logarithmic Sobolev equation. Using Herbst’s argument, e.g. along the lines of Ledoux (1999), a standard consequence of the logarithmic Sobolev equation is that the distribution is sub-Gaussian with a parameter antiproportional to the parameter of the bound of the Hessian. For $t>t^{*}$, we have that
$\nabla^{2}\Psi(y)=\frac{1}{s^{2}}I+\nabla^{2}V(y)\gtrsim\Big(\frac{1}{s^{2}}-L\Big)I\gtrsim\frac{1}{2s^{2}}I.$
Note that we used a different property of $V$, namely the bound on the spectral norm of $\nabla^{2}V$. From the sub-Gaussianity, we can control the moment $\mathbb{E}[|Y-\mathbb{E}[Y]|^{3}]\lesssim s^{3}$, see Jin et al. (2019, Lemma 1). Since $\nabla^{2}V(\mathbb{E}[Y])$ is invertible, we obtain
$\Sigma=s^{2}(I-R^{\prime})(I+s^{2}(\nabla V(\mathbb{E}[Y]))^{\top})^{-1}.$
Next, we use a Neumann series, which is admissible since $\|s^{2}(\nabla V(\mathbb{E}[Y]))^{\top}\|\leq s^{2}L\leq s\leq 1$ to rewrite
$(I+s^{2}(\nabla V(\mathbb{E}[Y]))^{\top})^{-1}=I-s^{2}(\nabla V(\mathbb{E}[Y]))^{\top}+\sum_{k=2}^{\infty}(-s^{2}(\nabla V(\mathbb{E}[Y]))^{\top})^{k}.$
For the sum we have using a geometric series
$\|\sum_{k=2}^{\infty}(-s^{2}(\nabla V(\mathbb{E}[Y]))^{\top})^{k}\|\leq\sum_{k=2}^{\infty}\|s^{2}(\nabla V(\mathbb{E}[Y]))^{\top})\|^{k}\leq\sum_{k=2}^{\infty}s^{k}=\frac{s^{2}}{1-s}\leq s^{2}.$
Thus
$\Sigma_{ij}=s^{2}(\mathds{1}_{i=j}+O(s^{2})).$
Property (III), $\mathbb{P}^{*}$ of form (13):
@@
$q\propto\exp\Big(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}\Big)=\exp\Big(-\frac{1}{2}\Big(1+\frac{t^{2}}{\sigma_{t}^{2}}\Big)\Big(|y|^{2}-\Big\langle\frac{t}{\sigma_{t}^{2}}\frac{x}{1+\frac{t^{2}}{\sigma_{t}^{2}}},y\Big\rangle\Big)\Big)$ (21)
is a Gaussian distribution with variance $\big(1+\frac{t^{2}}{\sigma_{t}^{2}}\big)^{-1}I_{d}$. Hence the density defined above satisfies the Poincaré inequality with constant $\big(1+\frac{t^{2}}{\sigma_{t}^{2}}\big)$. Using the Holley-Strooke perturbation principle Holley and Stroock (1987), in the form of Ledoux (2001, Lemma 1.2), we can bound the Poincaré constant $\rho$ of the perturbed Gaussian via $\rho\geq e^{-4L}\big(1+\frac{t^{2}}{\sigma_{t}^{2}}\big).$ Thus
@@
we conclude that for all $t\in[0,1]$ $\operatorname{Cov}(Y^{x,t}_{i},(Y^{x,t}_{j})\leq e^{4L}.$ Hence we can set $C=e^{4L}$.
Property (II), $\mathbb{P}^{*}$ of form (13):
@@
$\mathbb{E}[\big(Y^{x,t}_{i}\big)^{2}]$ $=\frac{\int\big(\frac{x_{i}-z_{i}\sigma_{t}}{t}\big)^{2}\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}{\int\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}\big(\frac{x_{i}}{t}\big)^{2}-\frac{2x_{i}\sigma_{t}}{t^{2}}\frac{\int z_{i}\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}{\int\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}+\big(\frac{\sigma_{t}}{t}\big)^{2}\frac{\int z_{i}^{2}\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}{\int\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}.$
@@
$\mathbb{E}[(Y^{x,t}_{i})^{2}]-\mathbb{E}[Y^{x,t}_{i}]^{2}$ $=\Big(\frac{x_{i}}{t}\Big)^{2}-\frac{2x_{i}\sigma_{t}}{t^{2}}A(z_{i})+\Big(\frac{\sigma_{t}}{t}\Big)^{2}A(z_{i}^{2})-\Big(\frac{x_{i}}{t}-\frac{\sigma_{t}}{t}A(z_{i})\Big)^{2}\Big(\frac{\sigma_{t}}{t}\Big)^{2}\Big(A(z_{i}^{2})-A(z_{i})^{2}\Big).$ (22)
@@
$p_{Z}(z)$ $=\frac{\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})}{\int\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}z}=\frac{\exp\big(-\frac{|z|^{2}}{2}-\frac{|x-z\sigma_{t}|^{2}}{2t^{2}}-a(\frac{x-z\sigma_{t}}{t})\big)}{\int\exp\big(-\frac{|z|^{2}}{2}-\frac{|x-z\sigma_{t}|^{2}}{2t^{2}}-a(\frac{x-z\sigma_{t}}{t})\big)\;\mathrm{d}z}$
@@
$\mathbb{E}_{p_{Z}}[\partial_{z_{i}}\log(p(Z))]=\int\partial_{z_{i}}p(z)\;\mathrm{d}z=\int...\int_{z_{i}\in\mathbb{R}}\partial_{z_{i}}p(z)\;\mathrm{d}z_{i}\;\mathrm{d}(z_{1},...,z_{i-1},z_{z+1},...,z_{d})=0.$
@@
$0=-\mathbb{E}\Big[\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)\Big(Z_{i}-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big)-\frac{\sigma_{t}}{t}\partial_{z_{i}}a\Big(\frac{x-Z\sigma_{t}}{t}\Big)\Big]$
$\Longleftrightarrow\quad$ $\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)\Big(\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big)=\mathbb{E}\Big[\frac{\sigma_{t}}{t}\partial_{z_{i}}a\Big(\frac{x-Z\sigma_{t}}{t}\Big)\Big]$
@@
where $L$ is from (13). From (23) we can see that this bounds the influence of the perturbation function $a$ on the expected value. Now we bound the variance. To do so, we first note that
@@
$\mathbb{E}_{p_{Z}}[\partial^{2}_{z_{i}}\log(p_{Z}(Z))]=\Big(1+\Big(\frac{\sigma_{t}}{t}\Big)^{2}\Big)+\Big(\frac{\sigma_{t}}{t}\Big)^{2}\mathbb{E}_{p_{Z}}\Big[\partial^{2}_{z_{i}}a(\frac{x-\sigma_{t}z}{t})\Big].$
@@
$\mathbb{E}_{p_{Z}}[(\partial_{z_{i}}\log(p(Z)))^{2}]=\mathbb{E}_{p_{Z}}\Big[\Big(\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)\Big(Z_{i}-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big)+\frac{\sigma_{t}}{t}\partial_{z_{i}}a\Big(\frac{x-\sigma_{t}Z}{t}\Big)\Big)^{2}\Big]$
$=\mathbb{E}_{p_{Z}}\Big[\Big(\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)\Big(Z_{i}-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big)\Big)^{2}\Big]+2\mathbb{E}_{p_{Z}}\Big[\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)\Big(Z_{i}-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big)\frac{\sigma_{t}}{t}\partial_{z_{i}}a\Big(\frac{x-\sigma_{t}Z}{t}\Big)\Big]$
@@
$\mathbb{E}_{p_{Z}}\Big[\Big(\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)\Big(Z_{i}-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big)\Big)^{2}\Big]=\mathbb{E}_{p_{Z}}\Big[\Big(\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)\Big(Z_{i}-\mathbb{E}[Z_{i}]+\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big)\Big)^{2}\Big]$
$=\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)^{2}\mathbb{E}_{p_{Z}}\big[(Z_{i}-\mathbb{E}[Z_{i}])^{2}\big]+2\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)^{2}\Big(\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big)\mathbb{E}_{p_{Z}}\big[(Z_{i}-\mathbb{E}[Z_{i}])\big]$
$+\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)^{2}\Big(\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big)^{2}=\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)^{2}\Big(\mathbb{E}_{p_{Z}}\big[(Z_{i}-\mathbb{E}[Z_{i}])^{2}\big]+\Big(\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big)^{2}\Big).$
@@
$\mathbb{E}_{p_{Z}}\Big[\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)\Big(Z_{i}-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big)\frac{\sigma_{t}}{t}\partial_{z_{i}}a\Big(\frac{x-\sigma_{t}Z}{t}\Big)\Big]$
$=\mathbb{E}_{p_{Z}}\Big[\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)\Big(Z_{i}-\mathbb{E}[Z_{i}]+\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big)\frac{\sigma_{t}}{t}\partial_{z_{i}}a\Big(\frac{x-\sigma_{t}Z}{t}\Big)\Big]$
$=\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)\frac{\sigma_{t}}{t}\mathbb{E}_{p_{Z}}\Big[(Z_{i}-\mathbb{E}[Z_{i}])\partial_{z_{i}}a\Big(\frac{x-\sigma_{t}Z}{t}\Big)\Big]+\Big(1+\frac{\sigma_{t}^{2}}{t^{2}}\Big)\frac{\sigma_{t}}{t}\Big(\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big)\mathbb{E}_{p_{Z}}\Big[\partial_{z_{i}}a\Big(\frac{x-\sigma_{t}Z}{t}\Big)\Big].$